OpenStudy (anonymous):

On a 246 km trip, a guy wants to achieve an average speed of 64.6 km/h. if she drives the first half of the trip at an average of 48.2 km/h, what average speed must she have for the second half of the trip?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (kropot72):

The required total time for the complete trip is \[t _{total}=\frac{distance}{speed}=\frac{246}{64.6}\ hours\] Time taken for the first half of the trip is \[t _{1}=\frac{123}{48.2}\ hours\] The time needed for the second half of the trip is given by \[t _{2}=t _{total}-t _{1}=\frac{246}{64.6}-\frac{123}{48.2}=3.81-2.55=1.26\ hours\] Therefore the average speed needed for the second half of the trip is \[Required\ speed=\frac{123}{1.26}\ km/h\]